# Dynamic Parameter Memory: Temporary LoRA-Enhanced LLM for Long-Sequence Emotion Recognition in Conversation

### v1

@@ 4.3.2 Comparison Study
We compared our model with recent models for ERC in Table 2. To ensure fairness, all compared models used IEMOCAP’s conventional four-class method. For multimodal methods, we selected results from their audio modality. MERITS-L (Dutta and Ganapathy, 2025) leverages LLMs to assist with the ERC task, while other methods improve on traditional classifier approaches. The performance of SLLM-based emotion recognition generally surpassed traditional classifier approaches, confirming our conclusions from the ablation study. With the same experimental settings, our method achieved SOTA results. This demonstrates the effectiveness of our designed emotion SLLM combined with DPM inference for capturing emotions in long audio sequences.
4.3.3 Does Truncated Context Affect LLM’s Emotion Understanding?
@@
Although this assumption is not critical for verifying DPM’s effectiveness, it is important for applying LLM to ERC. Therefore, we conducted a comparative experiment.
@@
Table 3: Performance comparison between models with different context window lengths.
@@
5 Discussion
Although in this paper we only evaluate the performance of the DPM on the ERC task, in theory, DPM can be applied to a wide range of audio understanding tasks, such as intent recognition, automatic speech recognition, and speech conversation.
First, DPM has a structural advantage over conventional autoregressive inference methods in that it can handle infinitely long audio sequences. Second, even when processing audio of the same length, results on the ERC task demonstrate that DPM significantly enhances understanding performance, especially on ultra-long audio inputs.
In future work, we plan to explore the effectiveness of DPM across a broader range of audio tasks.
6 Conclusions
The inherently high frame rate in the speech modality severely limits the emotion understanding capabilities of SLLM. This paper proposes a Dynamic Parameter Memory (DPM) mechanism with contextual semantics and sentence-level emotion encoding, enabling processing of unlimited-length audio with limited context windows in SLLM. Our method progressively encodes sentence-level information and emotions into a temporary LoRA module during inference, effectively "memorizing" contextual information. We trained an emotion SLLM as the basis for DPM inference. We validated our method on ERC. Experimental results on the IEMOCAP dataset show that DPM significantly improves the emotion recognition capabilities of SLLM when processing long audio sequences, achieving state-of-the-art performance.
Limitations
While DPM successfully enables understanding of unlimited-length audio sequences with linear computational complexity, some limitations remain:
Computational Efficiency vs. Performance Trade-off.
Although DPM demonstrates significant advantages in processing long audio sequences, it comes at the cost of reduced computational efficiency. In our implementation, the number of forward and backward passes required for inference equals the number of sentences in the audio sample. In contrast, traditional autoregressive inference requires only a single forward pass. Without considering DPM’s advantages in long audio understanding, there exists a clear trade-off between computational efficiency and computational cost.
Unexplored Impact of Different Stepping Strategies.
This paper does not examine the performance differences between various stepping strategies in DPM inference. Our current implementation employs an intuitive sentence-by-sentence progression, where each forward and backward pass during inference advances one sentence at a time, aligning with human reading habits. However, this approach is relatively inefficient. Using fixed-stride stepping would undoubtedly improve DPM’s efficiency—for example, setting a stride of 1024 would process approximately 40 seconds of audio per update at a 25Hz feature frame rate. The experimental effects of these different stepping strategies remain to be explored further.
Privacy Concerns.
Beyond technical limitations, deploying Speech Emotion Recognition (SER) technologies like DPM, which can infer emotions from history audio sequences, raises privacy concerns. The primary risk involves potential misuse for surveillance or emotional profiling if deployed without individuals’ consent. Continuously analyzing emotions from lengthy conversations could lead to detailed emotional histories; improper access to such histories would be a severe breach of personal privacy. Therefore, future work and any deployment must prioritize robust data protection, clear user consent mechanisms, and strict guidelines to prevent the unauthorized use of sensitive emotional data extracted from audio.

### v2

@@ 4.3.2 Comparison Study
We compared our model with recent models for ERC in Table 2. To ensure fairness, all compared models used IEMOCAP’s conventional four-class method. For multimodal methods, we selected results from their audio modality. MERITS-L (Dutta & Ganapathy, 2025) leverages LLMs to assist with the ERC task, while other methods improve on traditional classifier approaches. The performance of SLLM-based emotion recognition generally surpassed traditional classifier approaches, confirming our conclusions from the ablation study. With the same experimental settings, our method achieved SOTA results on both the IEMOCAP and MELD datasets. This demonstrates the effectiveness of our designed emotion SLLM combined with DPM inference for capturing emotions in long audio sequences.
5 Conclusions
The inherently high frame rate in the speech modality severely limits the emotion understanding capabilities of SLLM. This paper proposes a Dynamic Parameter Memory (DPM) mechanism with contextual semantics and sentence-level emotion encoding, enabling processing of unlimited-length audio with limited context windows in SLLM. Our method progressively encodes sentence-level information and emotions into a temporary LoRA module during inference, effectively ”memorizing” contextual information. We trained an emotion SLLM as the basis for DPM inference. We validated our method on ERC. Experimental results on the IEMOCAP dataset show that DPM significantly improves the emotion recognition capabilities of SLLM when processing long audio sequences, achieving state-of-the-art performance.
Appendix A Appendix
A.1 Analysis of DPM Stepping Strategies
Our primary implementation of DPM employs a sentence-by-sentence progression, which, while intuitive, may not be the most efficient. To investigate this further, we conducted an additional experiment on the MELD dataset to explore the trade-off between performance and efficiency under different stepping strategies. We compare our original sentence-based approach against a fixed-stride stepping strategy during DPM inference. The training process remains identical, only the mechanism for updating the temporary LoRA module during inference is changed from advancing by one semantic sentence to advancing by a fixed number of tokens. We tested stride lengths of 64, 128, 256, 512, and 1024 tokens.
Table 3: Impact of DPM stepping strategies on performance on the MELD dataset.
Stepping Strategy WF1 (%)
Sentence-by-sentence (Baseline) 51.22
Fixed-Stride (64 tokens) 49.61
Fixed-Stride (128 tokens) 49.64
Fixed-Stride (256 tokens) 47.63
Fixed-Stride (512 tokens) 43.25
Fixed-Stride (1024 tokens) 37.87
The results, presented in Table 3, reveal several key insights. Firstly, the sentence-by-sentence strategy outperforms all fixed-stride settings. This suggests that semantic boundaries, such as sentences, provide more meaningful and coherent chunks of information for the DPM to memorize than arbitrary fixed-length segments. This aligns with the intuition that emotions are often expressed and evolve at the sentence level.
Secondly, a clear trend of performance degradation is observed as the fixed stride becomes more aggressive. The best performance among fixed-stride settings is achieved with a stride of 128 tokens (49.64% WF1), after which performance drops significantly. This indicates that there is an optimal ”glimpse” length for the DPM mechanism. Interestingly, the average utterance duration in the MELD dataset is 3.22 seconds, which corresponds to approximately 80 tokens at a 25Hz feature frame rate. The fact that the best-performing fixed stride (128 tokens) is in a similar range further reinforces the hypothesis that a sentence-like unit is the most effective quantum for emotional context updates.
Finally, we analyze the efficiency trade-off. While the sentence-by-sentence approach is most accurate, fixed-stride stepping offers substantial efficiency gains. For instance, a stride of 1024 tokens improves inference speed by a factor of 12.72 compared to the sentence-by-sentence approach, but this comes at a significant cost, with the WF1 score dropping by 13.35% (from 51.22% to 37.87%). A more moderate stride of 512 tokens improves inference speed by a factor of 6.26, while performance decreases by 7.97%. This analysis highlights a clear trade-off, allowing practitioners to choose a strategy that balances the need for accuracy with computational constraints.
A.2 Does Truncated Context Affect LLM’s Emotion Understanding?
@@
Table 4: Performance comparison between models with different context window lengths on the IEMOCAP dataset.